\providecommand{\href}[2]{#2}\begingroup\raggedright\begin{thebibliography}{10}

\bibitem{Gale:2013da}
C.~Gale, S.~Jeon and B.~Schenke, \emph{{Hydrodynamic Modeling of Heavy-Ion
  Collisions}}, \href{https://doi.org/10.1142/S0217751X13400113}{\emph{Int. J.
  Mod. Phys.} {\bfseries A28} (2013) 1340011}
  [\href{https://arxiv.org/abs/1301.5893}{{\ttfamily 1301.5893}}].

\bibitem{Bozek:2016yoj}
P.~Bozek, \emph{{Transverse-momentum–flow correlations in relativistic
  heavy-ion collisions}},
  \href{https://doi.org/10.1103/PhysRevC.93.044908}{\emph{Phys. Rev.}
  {\bfseries C93} (2016) 044908}
  [\href{https://arxiv.org/abs/1601.04513}{{\ttfamily 1601.04513}}].

\bibitem{Aad:2019fgl}
{\scshape ATLAS} collaboration, \emph{{Measurement of flow harmonics
  correlations with mean transverse momentum in lead-lead and proton-lead
  collisions at $\sqrt{s_{NN}}=5.02$ TeV with the ATLAS detector}},
  \href{https://doi.org/10.1140/epjc/s10052-019-7489-6}{\emph{Eur. Phys. J.}
  {\bfseries C79} (2019) 985}
  [\href{https://arxiv.org/abs/1907.05176}{{\ttfamily 1907.05176}}].

\bibitem{Schenke:2020uqq}
B.~Schenke, C.~Shen and D.~Teaney, \emph{{Transverse momentum fluctuations and
  their correlation with elliptic flow in nuclear collisions}},
  \href{https://arxiv.org/abs/2004.00690}{{\ttfamily 2004.00690}}.

\bibitem{Giacalone:2020byk}
G.~Giacalone, B.~Schenke and C.~Shen, \emph{{Observable signatures of initial
  state momentum anisotropies in nuclear collisions}},
  \href{https://arxiv.org/abs/2006.15721}{{\ttfamily 2006.15721}}.

\bibitem{Carzon:2020ohg}
P.~Carzon, M.D.~Sievert and J.~Noronha-Hostler, \emph{{Importance of
  Multiplicity Fluctuations in Entropy Scaling}},  2020
  [\href{https://arxiv.org/abs/2007.12977}{{\ttfamily 2007.12977}}].

\bibitem{Moreland:2014oya}
J.S.~Moreland, J.E.~Bernhard and S.A.~Bass, \emph{{Alternative ansatz to
  wounded nucleon and binary collision scaling in high-energy nuclear
  collisions}}, \href{https://doi.org/10.1103/PhysRevC.92.011901}{\emph{Phys.
  Rev.} {\bfseries C92} (2015) 011901}.

\bibitem{McLerran:1993ni}
L.D.~McLerran and R.~Venugopalan, \emph{{Computing quark and gluon distribution
  functions for very large nuclei}},
  \href{https://doi.org/10.1103/PhysRevD.49.2233}{\emph{Phys. Rev.} {\bfseries
  D49} (1994) 2233} [\href{https://arxiv.org/abs/hep-ph/9309289}{{\ttfamily
  hep-ph/9309289}}].

\bibitem{Romatschke:2017ejr}
P.~Romatschke and U.~Romatschke, \emph{{Relativistic Fluid Dynamics In and Out
  of Equilibrium}}, Cambridge Monographs on Mathematical Physics, Cambridge
  University Press (5, 2019),
  \href{https://doi.org/10.1017/9781108651998}{10.1017/9781108651998},
  [\href{https://arxiv.org/abs/1712.05815}{{\ttfamily 1712.05815}}].

\bibitem{Chatrchyan:2012ta}
{\scshape CMS} collaboration, \emph{{Measurement of the elliptic anisotropy of
  charged particles produced in PbPb collisions at $\sqrt{s}_{NN}$=2.76 TeV}},
  \href{https://doi.org/10.1103/PhysRevC.87.014902}{\emph{Phys. Rev.}
  {\bfseries C87} (2013) 014902}
  [\href{https://arxiv.org/abs/1204.1409}{{\ttfamily 1204.1409}}].

\bibitem{Adam:2016ows}
{\scshape ALICE} collaboration, \emph{{Pseudorapidity dependence of the
  anisotropic flow of charged particles in Pb-Pb collisions at $\sqrt{s_{\rm
  NN}}=2.76$ TeV}},
  \href{https://doi.org/10.1016/j.physletb.2016.07.017}{\emph{Phys. Lett.}
  {\bfseries B762} (2016) 376}
  [\href{https://arxiv.org/abs/1605.02035}{{\ttfamily 1605.02035}}].

\bibitem{Aad:2020gfz}
{\scshape ATLAS} collaboration, \emph{{Longitudinal flow decorrelations in
  Xe+Xe collisions at $\sqrt{s_{\mathrm{NN}}}=5.44$ TeV with the ATLAS
  detector}},  \href{https://arxiv.org/abs/2001.04201}{{\ttfamily 2001.04201}}.

\bibitem{Shen:2017bsr}
C.~Shen and B.~Schenke, \emph{{Dynamical initial state model for relativistic
  heavy-ion collisions}},
  \href{https://doi.org/10.1103/PhysRevC.97.024907}{\emph{Phys. Rev.}
  {\bfseries C97} (2018) 024907}
  [\href{https://arxiv.org/abs/1710.00881}{{\ttfamily 1710.00881}}].

\bibitem{Shen:2020jwv}
C.~Shen and S.~Alzhrani, \emph{{A collision geometry-based 3D initial condition
  for relativistic heavy-ion collisions}},
  \href{https://doi.org/10.1103/PhysRevC.102.014909}{\emph{Phys. Rev.}
  {\bfseries C102} (2020) 014909}
  [\href{https://arxiv.org/abs/2003.05852}{{\ttfamily 2003.05852}}].

\bibitem{Mantysaari:2016ykx}
H.~Mäntysaari and B.~Schenke, \emph{{Evidence of strong proton shape
  fluctuations from incoherent diffraction}},
  \href{https://doi.org/10.1103/PhysRevLett.117.052301}{\emph{Phys. Rev. Lett.}
  {\bfseries 117} (2016) 052301}
  [\href{https://arxiv.org/abs/1603.04349}{{\ttfamily 1603.04349}}].

\bibitem{Tu:2020}
{\scshape STAR} collaboration in \emph{these proceedings}, 2020.

\bibitem{Mantysaari:2019jhh}
H.~Mäntysaari and B.~Schenke, \emph{{Accessing the gluonic structure of light
  nuclei at a future electron-ion collider}},
  \href{https://doi.org/10.1103/PhysRevC.101.015203}{\emph{Phys. Rev. C}
  {\bfseries 101} (2020) 015203}
  [\href{https://arxiv.org/abs/1910.03297}{{\ttfamily 1910.03297}}].

\bibitem{Tu:2020ymk}
Z.~Tu, A.~Jentsch, M.~Baker, L.~Zheng, J.-H.~Lee, R.~Venugopalan et~al.,
  \emph{{Probing short-range correlations in the deuteron via incoherent
  diffractive $J/\psi$ production with spectator tagging at the EIC}},
  \href{https://arxiv.org/abs/2005.14706}{{\ttfamily 2005.14706}}.

\bibitem{Cacciari:2012ny}
M.~Cacciari, S.~Frixione, N.~Houdeau, M.L.~Mangano, P.~Nason and G.~Ridolfi,
  \emph{{Theoretical predictions for charm and bottom production at the LHC}},
  \href{https://doi.org/10.1007/JHEP10(2012)137}{\emph{JHEP} {\bfseries 10}
  (2012) 137} [\href{https://arxiv.org/abs/1205.6344}{{\ttfamily 1205.6344}}].

\bibitem{Ma:2014mri}
Y.-Q.~Ma and R.~Venugopalan, \emph{{Comprehensive Description of J/$\psi$
  Production in Proton-Proton Collisions at Collider Energies}},
  \href{https://doi.org/10.1103/PhysRevLett.113.192301}{\emph{Phys. Rev. Lett.}
  {\bfseries 113} (2014) 192301}
  [\href{https://arxiv.org/abs/1408.4075}{{\ttfamily 1408.4075}}].

\bibitem{Acharya:2018kxc}
{\scshape ALICE} collaboration, \emph{{Inclusive J/$\psi$ production at forward
  and backward rapidity in p-Pb collisions at $\sqrt{s_{\rm NN}}$ = 8.16 TeV}},
  \href{https://doi.org/10.1007/JHEP07(2018)160}{\emph{JHEP} {\bfseries 07}
  (2018) 160} [\href{https://arxiv.org/abs/1805.04381}{{\ttfamily
  1805.04381}}].

\bibitem{Acharya:2020wwy}
{\scshape ALICE} collaboration, \emph{{Measurement of nuclear effects on
  $\psi\rm{(2S)}$ production in p-Pb collisions at $\sqrt{\textit{s}_{\rm NN}}
  = 8.16$ TeV}},  \href{https://arxiv.org/abs/2003.06053}{{\ttfamily
  2003.06053}}.

\bibitem{Ducloue:2016pqr}
B.~Ducloué, T.~Lappi and H.~Mäntysaari, \emph{{Forward $J/\psi$ production at
  high energy: centrality dependence and mean transverse momentum}},
  \href{https://doi.org/10.1103/PhysRevD.94.074031}{\emph{Phys. Rev.}
  {\bfseries D94} (2016) 074031}
  [\href{https://arxiv.org/abs/1605.05680}{{\ttfamily 1605.05680}}].

\bibitem{Albacete:2017qng}
J.L.~Albacete et~al., \emph{{Predictions for Cold Nuclear Matter Effects in
  $p+$Pb Collisions at $\sqrt{s_{_{NN}}} = 8.16$ TeV}},
  \href{https://doi.org/10.1016/j.nuclphysa.2017.11.015}{\emph{Nucl. Phys.}
  {\bfseries A972} (2018) 18}
  [\href{https://arxiv.org/abs/1707.09973}{{\ttfamily 1707.09973}}].

\bibitem{Kusina:2017gkz}
A.~Kusina, J.-P.~Lansberg, I.~Schienbein and H.-S.~Shao, \emph{{Gluon Shadowing
  in Heavy-Flavor Production at the LHC}},
  \href{https://doi.org/10.1103/PhysRevLett.121.052004}{\emph{Phys. Rev. Lett.}
  {\bfseries 121} (2018) 052004}
  [\href{https://arxiv.org/abs/1712.07024}{{\ttfamily 1712.07024}}].

\bibitem{Arleo:2014oha}
F.~Arleo and S.~Peigné, \emph{{Quarkonium suppression in heavy-ion collisions
  from coherent energy loss in cold nuclear matter}},
  \href{https://doi.org/10.1007/JHEP10(2014)073}{\emph{JHEP} {\bfseries 10}
  (2014) 073} [\href{https://arxiv.org/abs/1407.5054}{{\ttfamily 1407.5054}}].

\bibitem{Acharya:2020giw}
{\scshape ALICE} collaboration, \emph{{J/$\psi$ production as a function of
  charged-particle multiplicity in p-Pb collisions at $\sqrt{\textit{s}_{\rm
  NN}}~=~8.16$ TeV}},  \href{https://arxiv.org/abs/2004.12673}{{\ttfamily
  2004.12673}}.

\bibitem{Acharya:2019zjt}
{\scshape PHENIX} collaboration, \emph{{Measurement of $J/\psi$ at forward and
  backward rapidity in $p+p$, $p+A$l, $p+A$u, and $^3$He$+$Au collisions at
  $\sqrt{s_{_{NN}}}=200~{\rm GeV}$}},
  \href{https://doi.org/10.1103/PhysRevC.102.014902}{\emph{Phys. Rev.}
  {\bfseries C102} (2020) 014902}.

\bibitem{Acharya:2020puh}
{\scshape ALICE} collaboration, \emph{{Z-boson production in p-Pb collisions at
  $\sqrt{s_{\mathrm{NN}}}=8.16$ TeV and Pb-Pb collisions at
  $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV}},
  \href{https://arxiv.org/abs/2005.11126}{{\ttfamily 2005.11126}}.

\bibitem{Eskola:2020lee}
K.J.~Eskola, I.~Helenius, M.~Kuha and H.~Paukkunen, \emph{{Shadowing in
  inelastic nucleon-nucleon cross section?}},
  \href{https://arxiv.org/abs/2003.11856}{{\ttfamily 2003.11856}}.

\bibitem{Aad:2019sfe}
{\scshape ATLAS} collaboration, \emph{{Measurement of $W^\pm $ boson production
  in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}} = 5.02~\text {Te}\text {V}$
  with the ATLAS detector}},
  \href{https://doi.org/10.1140/epjc/s10052-019-7439-3}{\emph{Eur. Phys. J.}
  {\bfseries C79} (2019) 935}
  [\href{https://arxiv.org/abs/1907.10414}{{\ttfamily 1907.10414}}].

\bibitem{Aad:2019lan}
{\scshape ATLAS} collaboration, \emph{{$Z$ boson production in Pb+Pb collisions
  at $\sqrt{s_{\textrm{NN}}}$= 5.02 TeV measured by the ATLAS experiment}},
  \href{https://doi.org/10.1016/j.physletb.2020.135262}{\emph{Phys. Lett.}
  {\bfseries B802} (2020) 135262}
  [\href{https://arxiv.org/abs/1910.13396}{{\ttfamily 1910.13396}}].

\bibitem{Lappi:2013zma}
T.~Lappi and H.~Mäntysaari, \emph{{Single inclusive particle production at
  high energy from HERA data to proton-nucleus collisions}},
  \href{https://doi.org/10.1103/PhysRevD.88.114020}{\emph{Phys. Rev.}
  {\bfseries D88} (2013) 114020}
  [\href{https://arxiv.org/abs/1309.6963}{{\ttfamily 1309.6963}}].

\bibitem{Arleo:2020hat}
F.~Arleo, F.~Cougoulic and S.~Peigné, \emph{{Fully coherent energy loss
  effects on light hadron production in pA collisions}},
  \href{https://arxiv.org/abs/2003.06337}{{\ttfamily 2003.06337}}.

\bibitem{Adler:2006wg}
{\scshape PHENIX} collaboration, \emph{{Centrality dependence of pi0 and eta
  production at large transverse momentum in s(NN)**(1/2) = 200-GeV d+Au
  collisions}},
  \href{https://doi.org/10.1103/PhysRevLett.98.172302}{\emph{Phys. Rev. Lett.}
  {\bfseries 98} (2007) 172302}.

\bibitem{McGlinchey:2016ssj}
D.~McGlinchey, J.L.~Nagle and D.V.~Perepelitsa, \emph{{Consequences of high-$x$
  proton size fluctuations in small collision systems at $\sqrt {s_{NN}}=$
  200GeV}}, \href{https://doi.org/10.1103/PhysRevC.94.024915}{\emph{Phys. Rev.}
  {\bfseries C94} (2016) 024915}.

\bibitem{Wei:2020}
G.~Giacalone, C.~Marquet, M.~Matas and S.~Wei in \emph{these proceedings},
  2020.

\bibitem{Sirunyan:2018qel}
{\scshape CMS} collaboration, \emph{{Constraining gluon distributions in nuclei
  using dijets in proton-proton and proton-lead collisions at
  $\sqrt{s_{_\mathrm{NN}}} =$ 5.02 TeV}},
  \href{https://doi.org/10.1103/PhysRevLett.121.062002}{\emph{Phys. Rev. Lett.}
  {\bfseries 121} (2018) 062002}.

\bibitem{Balitsky:2008zza}
I.~Balitsky and G.A.~Chirilli, \emph{{Next-to-leading order evolution of color
  dipoles}}, \href{https://doi.org/10.1103/PhysRevD.77.014019}{\emph{Phys.
  Rev.} {\bfseries D77} (2008) 014019}
  [\href{https://arxiv.org/abs/0710.4330}{{\ttfamily 0710.4330}}].

\bibitem{Iancu:2015vea}
E.~Iancu, J.D.~Madrigal, A.H.~Mueller, G.~Soyez and D.N.~Triantafyllopoulos,
  \emph{{Resumming double logarithms in the QCD evolution of color dipoles}},
  \href{https://doi.org/10.1016/j.physletb.2015.03.068}{\emph{Phys. Lett.}
  {\bfseries B744} (2015) 293}.

\bibitem{Lappi:2015fma}
T.~Lappi and H.~Mäntysaari, \emph{{Direct numerical solution of the coordinate
  space Balitsky-Kovchegov equation at next to leading order}},
  \href{https://doi.org/10.1103/PhysRevD.91.074016}{\emph{Phys. Rev.}
  {\bfseries D91} (2015) 074016}.

\bibitem{Lappi:2016fmu}
T.~Lappi and H.~Mäntysaari, \emph{{Next-to-leading order Balitsky-Kovchegov
  equation with resummation}},
  \href{https://doi.org/10.1103/PhysRevD.93.094004}{\emph{Phys. Rev.}
  {\bfseries D93} (2016) 094004}
  [\href{https://arxiv.org/abs/1601.06598}{{\ttfamily 1601.06598}}].

\bibitem{Beuf:2020dxl}
G.~Beuf, H.~Hänninen, T.~Lappi and H.~Mäntysaari, \emph{{Color Glass
  Condensate at next-to-leading order meets HERA data}},
  \href{https://arxiv.org/abs/2007.01645}{{\ttfamily 2007.01645}}.

\bibitem{JalilianMarian:1997dw}
J.~Jalilian-Marian, A.~Kovner and H.~Weigert, \emph{{The Wilson renormalization
  group for low x physics: Gluon evolution at finite parton density}},
  \href{https://doi.org/10.1103/PhysRevD.59.014015}{\emph{Phys. Rev.}
  {\bfseries D59} (1998) 014015}.

\bibitem{Iancu:2001ad}
E.~Iancu, A.~Leonidov and L.D.~McLerran, \emph{{The Renormalization group
  equation for the color glass condensate}},
  \href{https://doi.org/10.1016/S0370-2693(01)00524-X}{\emph{Phys. Lett.}
  {\bfseries B510} (2001) 133}
  [\href{https://arxiv.org/abs/hep-ph/0102009}{{\ttfamily hep-ph/0102009}}].

\bibitem{Balitsky:2013fea}
I.~Balitsky and G.A.~Chirilli, \emph{{Rapidity evolution of Wilson lines at the
  next-to-leading order}},
  \href{https://doi.org/10.1103/PhysRevD.88.111501}{\emph{Phys. Rev.}
  {\bfseries D88} (2013) 111501}
  [\href{https://arxiv.org/abs/1309.7644}{{\ttfamily 1309.7644}}].

\bibitem{Kovner:2013ona}
A.~Kovner, M.~Lublinsky and Y.~Mulian, \emph{{Jalilian-Marian, Iancu, McLerran,
  Weigert, Leonidov, Kovner evolution at next to leading order}},
  \href{https://doi.org/10.1103/PhysRevD.89.061704}{\emph{Phys. Rev.}
  {\bfseries D89} (2014) 061704}.

\bibitem{Rummukainen:2003ns}
K.~Rummukainen and H.~Weigert, \emph{{Universal features of JIMWLK and BK
  evolution at small x}},
  \href{https://doi.org/10.1016/j.nuclphysa.2004.03.219}{\emph{Nucl. Phys.}
  {\bfseries A739} (2004) 183}
  [\href{https://arxiv.org/abs/hep-ph/0309306}{{\ttfamily hep-ph/0309306}}].

\bibitem{Dumitru:2011vk}
A.~Dumitru, J.~Jalilian-Marian, T.~Lappi, B.~Schenke and R.~Venugopalan,
  \emph{{Renormalization group evolution of multi-gluon correlators in high
  energy QCD}},
  \href{https://doi.org/10.1016/j.physletb.2011.11.002}{\emph{Phys. Lett.}
  {\bfseries B706} (2011) 219}
  [\href{https://arxiv.org/abs/1108.4764}{{\ttfamily 1108.4764}}].

\bibitem{Lappi:2020srm}
T.~Lappi, H.~Mäntysaari and A.~Ramnath, \emph{{Next-to-leading order
  Balitsky-Kovchegov equation beyond large $N_\mathrm{c}$}},
  \href{https://arxiv.org/abs/2007.00751}{{\ttfamily 2007.00751}}.

\bibitem{Roy:2018jxq}
K.~Roy and R.~Venugopalan, \emph{{Inclusive prompt photon production in
  electron-nucleus scattering at small x}},
  \href{https://doi.org/10.1007/JHEP05(2018)013}{\emph{JHEP} {\bfseries 05}
  (2018) 013} [\href{https://arxiv.org/abs/1802.09550}{{\ttfamily
  1802.09550}}].

\bibitem{Roy:2019hwr}
K.~Roy and R.~Venugopalan, \emph{{NLO impact factor for inclusive
  photon$+$dijet production in $e+A$ DIS at small $x$}},
  \href{https://doi.org/10.1103/PhysRevD.101.034028}{\emph{Phys. Rev.}
  {\bfseries D101} (2020) 034028}
  [\href{https://arxiv.org/abs/1911.04530}{{\ttfamily 1911.04530}}].

\bibitem{Roy:2019cux}
K.~Roy and R.~Venugopalan, \emph{{Extracting many-body correlators of saturated
  gluons with precision from inclusive photon+dijet final states in deeply
  inelastic scattering}},
  \href{https://doi.org/10.1103/PhysRevD.101.071505}{\emph{Phys. Rev.}
  {\bfseries D101} (2020) 071505}
  [\href{https://arxiv.org/abs/1911.04519}{{\ttfamily 1911.04519}}].

\bibitem{Kolbe:2020tlq}
I.~Kolbé, K.~Roy, F.~Salazar, B.~Schenke and R.~Venugopalan, \emph{{Inclusive
  prompt photon-jet correlations as a probe of gluon saturation in
  electron-nucleus scattering at small $x$}},
  \href{https://arxiv.org/abs/2008.04372}{{\ttfamily 2008.04372}}.

\bibitem{Li:2020uhl}
M.~Li, X.~Zhao, P.~Maris, G.~Chen, Y.~Li, K.~Tuchin et~al.,
  \emph{{Ultrarelativistic quark-nucleus scattering in a light-front
  Hamiltonian approach}},
  \href{https://doi.org/10.1103/PhysRevD.101.076016}{\emph{Phys. Rev.}
  {\bfseries D101} (2020) 076016}
  [\href{https://arxiv.org/abs/2002.09757}{{\ttfamily 2002.09757}}].

\bibitem{Jalilian-Marian:2020dwf}
J.~Jalilian-Marian, \emph{{High energy scattering in QCD: from low to high
  Bjorken $x$}}, \href{https://doi.org/10.22323/1.374.0047}{\emph{PoS}
  {\bfseries LC2019} (2020) 047}.

\bibitem{Boussarie:2020vzf}
R.~Boussarie and Y.~Mehtar-Tani, \emph{{On gauge invariance of transverse
  momentum dependent distributions at small x}},
  \href{https://arxiv.org/abs/2001.06449}{{\ttfamily 2001.06449}}.

\bibitem{Duan:2020jkz}
H.~Duan, C.~Akkaya, A.~Kovner and V.V.~Skokov, \emph{{Entanglement, partial set
  of measurements, and diagonality of the density matrix in the parton model}},
  \href{https://doi.org/10.1103/PhysRevD.101.036017}{\emph{Phys. Rev.}
  {\bfseries D101} (2020) 036017}
  [\href{https://arxiv.org/abs/2001.01726}{{\ttfamily 2001.01726}}].

\bibitem{Florkowski:2017olj}
W.~Florkowski, M.P.~Heller and M.~Spalinski, \emph{{New theories of
  relativistic hydrodynamics in the LHC era}},
  \href{https://doi.org/10.1088/1361-6633/aaa091}{\emph{Rept. Prog. Phys.}
  {\bfseries 81} (2018) 046001}
  [\href{https://arxiv.org/abs/1707.02282}{{\ttfamily 1707.02282}}].

\bibitem{Berges:2020fwq}
J.~Berges, M.P.~Heller, A.~Mazeliauskas and R.~Venugopalan,
  \emph{{Thermalization in QCD: theoretical approaches, phenomenological
  applications, and interdisciplinary connections}},
  \href{https://arxiv.org/abs/2005.12299}{{\ttfamily 2005.12299}}.

\bibitem{Kurkela:2019set}
A.~Kurkela, W.~van~der Schee, U.A.~Wiedemann and B.~Wu, \emph{{Early- and
  Late-Time Behavior of Attractors in Heavy-Ion Collisions}},
  \href{https://doi.org/10.1103/PhysRevLett.124.102301}{\emph{Phys. Rev. Lett.}
  {\bfseries 124} (2020) 102301}
  [\href{https://arxiv.org/abs/1907.08101}{{\ttfamily 1907.08101}}].

\bibitem{Kurkela:2018xxd}
A.~Kurkela and A.~Mazeliauskas, \emph{{Chemical Equilibration in Hadronic
  Collisions}},
  \href{https://doi.org/10.1103/PhysRevLett.122.142301}{\emph{Phys. Rev. Lett.}
  {\bfseries 122} (2019) 142301}
  [\href{https://arxiv.org/abs/1811.03040}{{\ttfamily 1811.03040}}].

\bibitem{Kurkela:2018oqw}
A.~Kurkela and A.~Mazeliauskas, \emph{{Chemical equilibration in weakly coupled
  QCD}}, \href{https://doi.org/10.1103/PhysRevD.99.054018}{\emph{Phys. Rev.}
  {\bfseries D99} (2019) 054018}
  [\href{https://arxiv.org/abs/1811.03068}{{\ttfamily 1811.03068}}].

\bibitem{Du:2020}
X.~Du and S.~Schlichting in \emph{these proceedings}, 2020.

\bibitem{Gale:2020xlg}
C.~Gale, J.-F.~Paquet, B.~Schenke and C.~Shen, \emph{{Probing Early-Time
  Dynamics and Quark-Gluon Plasma Transport Properties with Photons and
  Hadrons}},  in \emph{{Proceedings of Quark Matter 2019, Wuhan, China,
  November 4-9, 2019}}, 2020
  [\href{https://arxiv.org/abs/2002.05191}{{\ttfamily 2002.05191}}].

\bibitem{Schenke:2012wb}
B.~Schenke, P.~Tribedy and R.~Venugopalan, \emph{{Fluctuating Glasma initial
  conditions and flow in heavy ion collisions}},
  \href{https://doi.org/10.1103/PhysRevLett.108.252301}{\emph{Phys. Rev. Lett.}
  {\bfseries 108} (2012) 252301}
  [\href{https://arxiv.org/abs/1202.6646}{{\ttfamily 1202.6646}}].

\bibitem{Kurkela:2018vqr}
A.~Kurkela, A.~Mazeliauskas, J.-F.~Paquet, S.~Schlichting and D.~Teaney,
  \emph{{Effective kinetic description of event-by-event pre-equilibrium
  dynamics in high-energy heavy-ion collisions}},
  \href{https://doi.org/10.1103/PhysRevC.99.034910}{\emph{Phys. Rev.}
  {\bfseries C99} (2019) 034910}
  [\href{https://arxiv.org/abs/1805.00961}{{\ttfamily 1805.00961}}].

\bibitem{Schenke:2010nt}
B.~Schenke, S.~Jeon and C.~Gale, \emph{{(3+1)D hydrodynamic simulation of
  relativistic heavy-ion collisions}},
  \href{https://doi.org/10.1103/PhysRevC.82.014903}{\emph{Phys. Rev.}
  {\bfseries C82} (2010) 014903}
  [\href{https://arxiv.org/abs/1004.1408}{{\ttfamily 1004.1408}}].

\bibitem{Schenke:2010rr}
B.~Schenke, S.~Jeon and C.~Gale, \emph{{Elliptic and triangular flow in
  event-by-event (3+1)D viscous hydrodynamics}},
  \href{https://doi.org/10.1103/PhysRevLett.106.042301}{\emph{Phys. Rev. Lett.}
  {\bfseries 106} (2011) 042301}
  [\href{https://arxiv.org/abs/1009.3244}{{\ttfamily 1009.3244}}].

\bibitem{Bass:1998ca}
S.A.~Bass et~al., \emph{{Microscopic models for ultrarelativistic heavy ion
  collisions}},
  \href{https://doi.org/10.1016/S0146-6410(98)00058-1}{\emph{Prog. Part. Nucl.
  Phys.} {\bfseries 41} (1998) 255}
  [\href{https://arxiv.org/abs/nucl-th/9803035}{{\ttfamily nucl-th/9803035}}].

\end{thebibliography}\endgroup

\documentclass[a4paper,11pt]{article}
\usepackage[utf8]{inputenc}
\usepackage{pos}
\usepackage{natbib}
\title{Initial State Summary of Hard Probes 2020}
\author*[a]{Bj\"orn Schenke}
\affiliation[a]{Physics Department, Brookhaven National Laboratory, Bldg. 510A, Upton, NY 11973, USA}

\emailAdd{bschenke@bnl.gov}

\abstract{The description of the initial state of heavy ion collisions, which covers the description of the incoming nuclei, the initial hard and soft interactions, the resulting spatial geometry of the produced matter, as well as the dynamic approach to a medium well described by hydrodynamics, has important consequences for the study of hard and electromagnetic probes. I will review new developments presented at Hard Probes 2020 that have an impact on these aspects of our understanding of the initial state of heavy ion and smaller system collisions.}

\FullConference{%
  HardProbes2020\\
  1-6 June 2020\\
  Austin, Texas}

\begin{document}
\maketitle

%\section{Introduction}
\section{Initial geometry}

It has been established that the azimuthal momentum anisotropy of produced hadrons in heavy ion collisions originates from the (hydrodynamic) response of the produced medium to the initial geometry in the plane transverse to the beamline \cite{Gale:2013da}. Understanding the precise geometry and its fluctuations over a wide range of collision systems and energies is thus crucial for describing the bulk observables in heavy ion collisions. 

A powerful new observable that correlates the elliptic anisotropy $v_2$ and the mean transverse momentum was suggested in \cite{Bozek:2016yoj}, and measurements by the ATLAS Collaboration \cite{Aad:2019fgl} were presented. The correlator is sensitive to the details of the initial state model and dominated by the initial geometry \cite{Bozek:2016yoj, Schenke:2020uqq} in all systems, as long as $dN_{\rm ch}/d\eta>10$ . For smaller multiplicities initial momentum anisotropies will begin to dominate the final correlator, resulting in a characteristic sign change as a function of multiplicity \cite{Giacalone:2020byk}. Experiments have begun to investigate this, but are challenged by the presence of ``non-flow'' contributions at small multiplicities.

Other observables that can distinguish different models, for example the difference between energy deposition $\sim \sqrt{T_A T_B}$ and $\sim T_A T_B$ (where $T_{A/B}$ are the position dependent target and projectile thickness functions), include the ratio of the four and two particle cumulants $v_n\{4\}/v_n\{2\}$ \cite{Carzon:2020ohg}. The prescription $\sim \sqrt{T_A T_B}$ is preferred by the Trento model \cite{Moreland:2014oya}, which does not include the possibility of $\sim T_A T_B$ scaling, which is predicted by the color glass condensate (CGC) effective field theory (EFT) / glasma \cite{McLerran:1993ni} %,McLerran:1993ka} 
and AdS/CFT \cite{Romatschke:2017ejr}. 

We particularly need to better understand the rapidity dependence of the initial state. Even at the highest LHC energies, measurements of flow harmonics show a rather strong rapidity dependence \cite{Chatrchyan:2012ta,Adam:2016ows}, especially in more peripheral events. More complex observables can provide better constraints for 3+1 dimensional models, for example the longitudinal flow decorrelation, which was studied in Xe+Xe and Pb+Pb systems by the ATLAS Collaboration \cite{Aad:2020gfz}. This measurement revealed that hydrodynamic models that can describe the ratios of harmonic flow coefficients between Xe+Xe and Pb+Pb collisions, fail to describe most of the rapidity dependence characterized by $v_n$ rapidity decorrelations. New initial state models for the three dimensional structure of heavy ion collisions are being developed \cite{Shen:2017bsr,Shen:2020jwv}, and especially at low collision energies dynamical energy deposition is required, resulting in complex three dimensional fluctuating initial energy and net baryon density distributions \cite{Shen:2017bsr}. Detailed data as that described above will be very useful for constraining such models.

Finally, input that helps constrain the transverse geometry in small system collisions, such as p/d+A collisions, can be obtained from measurements of photoproduction of vector mesons, as was done in \cite{Mantysaari:2016ykx} using data from diffractive $J/\psi$ production in e+p collisions at HERA. New data on $J/\psi$ photoproduction in d+Au collisions was presented by the STAR Collaboration \cite{Tu:2020}. The computation of the coherent and incoherent cross section in the CGC framework is sensitive to the average shape and fluctutions of the deuteron, and the data seems to be able to constrain both \cite{Mantysaari:2019jhh}, and will hopefully also provide more insight into short range nucleon correlations \cite{Tu:2020ymk}.

\vspace{-0.4cm}
\section{Nuclear effects}
\vspace{-0.4cm}
To properly describe the initial state of heavy ion collisions and understand for example the background to the medium modification of jets, nuclear modifications of parton distribution functions (PDFs) need to be well understood. A range of observables can be used to constrain nuclear PDFs, including quarkonium, electroweak boson, light meson, as well as dijet production. 

$J/\psi$ production in p+p collisions can be well described within a combined CGC+non-relativistic QCD (NRQCD) (+Fixed Order+Next-to-Leading Log (FONLL) \cite{Cacciari:2012ny}) framework \cite{Ma:2014mri}. The nuclear modification observed in p+Pb collisions \cite{Acharya:2018kxc,Acharya:2020wwy} is also well described within the CGC EFT \cite{Ducloue:2016pqr,Albacete:2017qng}, and also using constrained nuclear PDFs \cite{Kusina:2017gkz,Albacete:2017qng} and coherent energy loss \cite{Arleo:2014oha,Albacete:2017qng}. What has not yet been fully described theoretically is the correlation between $J/\psi$ production at backward and forward rapidities and midrapidity charged hadron production in p+Pb collisions \cite{Acharya:2020giw}. It will be interesting to see what physics aspects are needed to describe the observed trends.

The nuclear dependence of $J/\psi$ production as a function of rapidity was studied by PHENIX at RHIC \cite{Acharya:2019zjt}. As expected, in p+Au collisions one finds stronger shadowing in the forward direction and stronger nuclear absorption in the backward direction compared to p+Al collisions. For collisions with a Au target (p+Au and $^3$He+Au), only nuclear PDFs are not sufficient to describe the measured $R_{AB}$ ratios. These types of measurements can be very useful to understand the dependence on both nuclear projectile and target. 

Electroweak boson production is also sensitive to nuclear effects, as demonstrated by ALICE for Z \cite{Acharya:2020puh} and W$^{\pm}$ production (preliminary data presented at this conference), and new forward measurements can be used as additional input to global nuclear PDF fits. Electroweak boson data has also been suggested to be used to calibrate the Glauber model \cite{Eskola:2020lee}: Assuming that EPPS16 correctly describes nuclear effects on the production of W and Z bosons, using a reduced nucleon-nucleon inelastic cross section of $41.5^{+16.2}_{-12.0}\,{\rm mb}$ compared to the usual $70\pm 5\,{\rm mb}$ at $\sqrt{s}=5\,{\rm TeV}$ leads to much better agreement with the $R_{\rm PbPb}$ for W and Z bosons measured at LHC \cite{Aad:2019sfe,Aad:2019lan}, implying nuclear shadowing of the nucleon-nucleon cross section itself. 

The nuclear modification of light mesons has now been measured out to $p_T$ of up to 200 GeV in 8.16 TeV p+Pb collisions by ALICE and presented at this conference. The results are consistent with theory calculations within the CGC EFT \cite{Lappi:2013zma} and coherent energy loss \cite{Arleo:2020hat}. Further insight can be gained from small system scan collisions at RHIC, where the $\pi^0$ $R_{p/d/^3He+Au}$ has been measured by PHENIX (\cite{Adler:2006wg} and preliminary data). Similar to p+Pb collisions at LHC, the centrality dependence observed is consistent with fluctuating proton sizes \cite{McGlinchey:2016ssj}.

Finally, dijet measurements in p+p and p+A collisions provide additional information on nuclear effects. For example, predictions for forward dijet angular correlations were presented within the CGC framework \cite{Wei:2020}. Because of Sudakov effects, which depend on the kinematics and can mask saturation effects, the ideal range in dijet transverse momenta to observe saturation effects, was identified to be 5 to 8 GeV (at forward rapidities) in $\sqrt{s}=8.8\,{\rm TeV}$ collisions. The CMS Collaboration presented measurements of the ratio of rapidity dependent dijet production in p+Pb over p+p collisions \cite{Sirunyan:2018qel}. This provided first evidence that apart from suppression at small $x$, the gluon PDF at large $x$ in Pb is strongly suppressed with respect to the PDF in unbound nucleons.

\vspace{-0.4cm}
\section{Theory progress at small x}
\vspace{-0.4cm}
Steady progress is being made in extending calculations within the CGC EFT to next-to-leading order (NLO). First phenomenological applications using the full NLO impact factor and the NLO Balitsky-Kovchegov (BK) equation \cite{Balitsky:2008zza,Iancu:2015vea,Lappi:2015fma,Lappi:2016fmu} to determine the $x$ dependence of deep inelastic scattering (DIS) observables have become available \cite{Beuf:2020dxl}. While the calculation does only include terms that are enhanced by large transverse logarithms, three different formulations of these evolution equations (which resum some or all of the transverse logarithms) yield little difference. Like at leading order (LO), there seems to be a significant nonperturbative contribution to the structure function for light quarks. Extension of the NLO calculation of the impact factor for heavy quarks, which would improve this situation, is still outstanding.

The BK equation can be viewed as the large $N_c$ limit of the JIMWLK equation \cite{JalilianMarian:1997dw,Iancu:2001ad}, which has also been extended to NLO \cite{Balitsky:2013fea,Kovner:2013ona}. At LO, it is known that finite $N_c$ corrections are small and the BK equation approximates the dipole evolution well \cite{Rummukainen:2003ns,Dumitru:2011vk}. At NLO, appearing six-point functions factorize into dipole operators at large $N_c$. Using the Gaussian approximation for these higher point correlators, a closed finite-$N_c$ BK equation at NLO was obtained \cite{Lappi:2020srm}. Again, it was shown that finite $N_c$ corrections are smaller than the naively expected $1/N_c^2\sim\mathcal{O}(10\%)$.

A complete NLO calculation of inclusive dijet+photon production in e+p/A collisions within the CGC EFT has recently been outlined in \cite{Roy:2018jxq,Roy:2019hwr,Roy:2019cux}. This involves in particular the calculation of the NLO impact factors. The calculation is powerful, as it includes many NLO processes at small $x$, such as inclusive dijet, inclusive photon+jet, and inclusive photon production, as well as fully inclusive DIS. It further carries information on LO $q\bar{q}g$ and $q\bar{q}g+\gamma$ production. Combining these results for the NLO impact factor with next-to-leading-log (NLL) resummed evolution kernels, an accuracy of order $\alpha_s^3\ln(1/x)$ can be achieved. Based on this, explicit calculations for prompt photon+jet production can be done at NLO. First LO results for this process, studying the azimuthal photon-jet correlations, which carry information on saturation, have recently been published in \cite{Kolbe:2020tlq}.

Further progress reported included extensions beyond the eikonal approximation: 1) in the calculation of quark-nucleus scattering in a light-front Hamiltonian approach \cite{Li:2020uhl}, and 2) by including large $x$ gluons in the target, which also aims at developing a unified picture for particle production at small and large $x$ \cite{Jalilian-Marian:2020dwf}. A generalized connection between transverse momentum dependent PDFs (TMDs) and the CGC EFT, using a new approach to TMDs at small $x$ employing transverse gauge link operators \cite{Boussarie:2020vzf} was presented, as well as the first computation of entanglement and ``ignorance'' entropies within the CGC EFT, the latter reflecting our inability to perform a complete set of measurements (sensitive to off-diagonal elements of the density matrix) \cite{Duan:2020jkz}.

\vspace{-0.4cm}
\section{Approach to equilibrium}
\vspace{-0.4cm}
Another area where a lot of progress has been made in the last couple of years is the study of the early time non-equlibrium evolution that is expected to result in a system that is well approximated by hydrodynamics \cite{Florkowski:2017olj,Berges:2020fwq}. The question of when hydrodynamics is applicable can be studied in both strong and weak coupling limits \cite{Kurkela:2019set}. In the strong coupling limit, described within AdS/CFT, causality requires that hydrodynamization occurs on time scales of order $\sim 1/T$, where $T$ is the temperature, as this is the distance from the horizon to the boundary, where the field theory lives. In the weak coupling limit, described by the Boltzmann equation and a relaxation time approximation (RTA) for the collision term, hydrodynamization is found to occur within $2\tau_0$, where $\tau_0$ is the initial time. The two limits behave somewhat differently, as in the RTA solutions approach the attractor before any finite order hydrodynamic expansion agrees with the attractor, while in AdS/CFT 2nd order hydrodynamics is usually a good approximation by the time the solutions approach the attractor.

Also chemical equilibration has been studied using the Boltzmann equation for quarks and gluons including $2\leftrightarrow 2$ and $1\leftrightarrow 2$ processes \cite{Kurkela:2018xxd,Kurkela:2018oqw}. It was found that chemical equilibration occurs after hydrodynamization but before local thermalization. Interestingly, chemical equilibration occurs for all systems with $dN_{\rm ch}/d\eta \gtrsim 100$, independent of the system size. Furter studies \cite{Du:2020} have found that chemical and kinetic equilibration for an expanding QGP with non-zero charge density occurs on roughly the same time scale as for zero charge density systems. Delayed chemical equilibration (relative to ``hydrodynamization'') can have an important effect on photon production and will affect photon anisotropic flow, as was explicitly shown in \cite{Gale:2020xlg} using a hybrid calculation employing IP-Glasma initial state \cite{Schenke:2012wb}, KoMPoST pre-equilibrium evolution \cite{Kurkela:2018vqr}, \textsc{Music} hydrodynamics \cite{Schenke:2010nt,Schenke:2010rr}, and UrQMD hadronic afterbuner \cite{Bass:1998ca}.

\vspace{-0.4cm}
\section{Conclusions}
\vspace{-0.4cm}
Our understanding of the initial state in heavy ion collisions and the early time dynamics is continuously improving. A lot of progress is being made in the first principles understanding of the nuclear wave function and nuclear and saturation effects, specifically the CGC EFT is being extended to NLO and calculations for a variety of processes at NLO are emerging. Furthermore, the initial geometry and its fluctuations are being better constrained, including the longitudinal direction, and the approach to a system that is described by hydrodynamics is being understood increasingly well. Among many other subfields of high energy nuclear physics, the study of hard and electromagnetic probes in heavy ion collisions will benefit immensly from these developments.

\vspace{-0.4cm}
\section*{Acknowledgments}
\vspace{-0.4cm}
B.P.S. is supported under DOE Contract No. DE-SC0012704.
\vspace{-0.4cm}

\bibliographystyle{JHEP} 
\bibliography{hp}

\end{document}